OpenStudy (anonymous):

Multiply seven-eighths times twenty-eight. thirty-two 4 and three-eighths one thirty-second twenty-four and one-half

OpenStudy (psymon):

Well, to help visually, we can first say that 28 is divided by one. So I'll write the multiplication like this: \[\frac{ 7 }{ 8 }*\frac{ 28 }{ 1 }\] Now when you multiply fractions, you simply multiply the two top numbers and then multiply the two bottom numbers. So this is what you will be doing: \[\frac{ 7*28 }{ 8*1 }\] That kind of make sense so far?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

7 x 28 is 196 Though. And is 8.

OpenStudy (anonymous):

It is?

OpenStudy (psymon):

I'm just going by what the question appeared to say O.o

OpenStudy (psymon):

But yes, 7*28 would be 196 for the top and the bottom would be 8. Now we would need to reduce it as much as possible from there.

OpenStudy (anonymous):

but 20 x 8 is 160. Which is none of the options.

OpenStudy (anonymous):

the answer is 24.5

OpenStudy (anonymous):

.... Thats not even an Option.

OpenStudy (psymon):

Anyone can just spit out an answer, but that's not the point, we want people to be able to figure it out on their own.

OpenStudy (psymon):

If you just want the answer then we can stop, but do you want to keep working through the problem or no?

OpenStudy (anonymous):

24.5 is twenty-four and one-half

OpenStudy (anonymous):

o.o

OpenStudy (psymon):

Some people just want answers given to them. So you can take that answer and we can stop or we can just work through the problem the rest of the way, your choice. I'm not here to make people listen to steps and work it out, I understand some people could care less about that and just want the answers. If so, then I'm sure soda will be glad to just give the rest of the answers to your problems.